22) Assume that losses fro m policyholders A, B, and C are independent and Expo- unt for policyholders A, B a d C oss d 700, respectively. What is the probability that the maximum is 200, 450, an loss from the three policyholders will exceed 900? A) 0.001 B) 0.381 C) 0.450 (D0.619 E) 1.000 The circled answer is wrong show clear work to arrive at a correct answer please.
Answer #1

P(maximum loss excceds 900) =P(max(A,B,C) >900)=1-P(max(A,B,C)<900)

=1-P(A<900)*P(B<900)*P(C<900)=1-(1-e-900/200)*(1-e-900/450)*(1-e-900/700)=0.381

option b